Two trains are 440 miles apart, and their speeds differ by 16 mph. Find the speed of each train if they are traveling toward each other and will meet in 4 hours.
Answer by macston(5194)

Let S=the speed of the slower train , S+16mph the speed of the faster train
The combined speed of the trains must be 440 mi/4hrs=110 mph
S+(S+16)=110 mph
2S+16=110mph subtract 16 from each side
2S=94 mph Divide each side by 2
S=47 mph ANSWER The slower train is traveling 47 mph
S+16=63mph ANSWER the faster train is traveling 63 mph
CHECK
In four hours, the slower train travels (4 hrs)(47 mph)= 188 miles
In four hours, the faster train travels (4 hrs)(63 mph)= 252 miles
188 miles + 252 miles=440 miles, which is the distance that separates them.